in a certain chemical reaction 1.00 g of product is produced in 15 minutes when the reaction is carried out at 20°C. Assuming that there is sufficient starting material available, how much product will be produced in the same length of time if the reaction is carried out in a boiling water bath at 100°C?
Answer #1

For every 10℃, rise in temperature , the rate of reaction increases by a factor of 2 (approximately) .

Number (n)of 10℃ rises =( 100-20)/10 = 8

Rate of the reaction at 100℃ = 28 × rate of reaction at 20℃

Rate of reaction at 100℃ = 256 × 1g/15mintues

Mass of product formed at 100℃ in 15 minutes = 256g/15minutes × 15 minutes = 256g

= 256g. (Answer)